OpenStudy (anonymous):

Find x in terms of a so that the line joining the points (x, x/a), (a, a/x) shall have a slope of 1

jimthompson5910 (jim_thompson5910):

Slope: m = (y2 - y1)/(x2 - x1) m = (a/x - x/a)/(a - x) m = (a^2/ax - x^2/ax)/(a - x) m = ((a^2 - x^2)/ax)/(a - x) m = (a^2 - x^2)/(ax(a - x)) m = ((a - x)(a + x))/(ax(a - x)) m = (a + x)/(ax) 1 = (a + x)/(ax) Now solve for x

OpenStudy (anonymous):

@jim_thompson5910 a=(x)/(x-1) ?

jimthompson5910 (jim_thompson5910):

you want to solve for x, not 'a'

OpenStudy (anonymous):

x=a/(a-1)?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks